# Agent-Oriented Planning in Multi-Agent Systems

### v1

@@ Abstract
Through the collaboration of multiple agents possessing diverse expertise and tools, multi-agent systems achieve impressive progress in solving real-world problems. Given the user queries, the meta-agents, serving as the brain within these systems, are required to decompose the queries into multiple sub-tasks that can be allocated to suitable agents capable of solving them, so-called agent-oriented planning. In this study, we identify three critical design principles of agent-oriented planning, including solvability, completeness, and non-redundancy, to ensure that each sub-task is effectively resolved, leading to satisfactory responses to the original queries. These principles further inspire us to propose a novel framework for agent-oriented planning in multi-agent systems, leveraging a fast task decomposition and allocation process followed by an effective and efficient evaluation via a reward model. During the planning process, the meta-agent is also responsible for evaluating the performance of the expert agents, making timely adjustments to the sub-tasks and scheduling as necessary. Besides, we integrate a feedback loop into the proposed framework to further enhance the effectiveness and robustness of such a problem-solving process. Extensive experiments demonstrate the advancement of the proposed framework in solving real-world problems compared to both single-agent systems and existing planning strategies for multi-agent systems.
@@ 1 Introduction
In recent years, large language models (LLMs) (Achiam et al., 2023; Chowdhery et al., 2023; Touvron et al., 2023) have achieved impressive breakthroughs in natural language understanding and generation, marking a critical advancement in the exploration of artificial general intelligence (AGI). As the capabilities of LLMs progress, LLM-empowered agents (Qin et al., 2023b; Dong et al., 2023; Wang et al., 2024) emerge as key components for integrating expertise and tools to translate these advancements into practical applications effectively. Advancing this paradigm, multi-agent systems (Shen et al., 2024; Chen et al., 2023; Wu et al., 2023b), which involve multiple diverse agents, provide great flexibility and adaptability by leveraging and collaborating on the strengths of various agents, promoting comprehensive solutions to complex real-world problems.
@@
Previous studies (Cai et al., 2023; Qian et al., 2023) propose to define suitable standard operating procedures (SOPs) based on the insights and experiences of human professionals for solving specific tasks, such as software development (Hong et al., 2023) and simulating personality traits (Serapio-García et al., 2023). In these scenarios, multiple agents are assigned to execute tasks following the predefined SOPs, resulting in remarkable successes. However, for multi-agent systems designed to tackle a diverse range of complex real-world problems, there is a critical need for a central entity that can automatically generate task-specific operating procedures and coordinate the activities of various agents (Wang et al., 2024).
@@
This central entity, typically referred to as a meta-agent (a.k.a. a controller or planner), carries two primary responsibilities. Firstly, the meta-agent needs to comprehend user queries and decompose them into several sub-tasks, ensuring that each sub-task can be adequately addressed by a single agent. Secondly, the meta-agent is also expected to assign these sub-tasks to the appropriate agents for execution, enabling the solutions of these sub-tasks to collectively provide comprehensive answers to the original user query.
@@
However, since the meta-agent cannot effectively associate sub-tasks based on the provided agent descriptions, the performance of task decomposition and allocations might be sub-optional. To tackle these challenges, we identify three design principles for an agent-oriented planning framework, including solvability, completeness, and non-redundancy, which further inspire us to propose a novel framework for multi-agent systems.
@@
Specifically, to resolve a user query, the meta-agent first performs fast task decomposition and allocation, which serves as the intermediate results that can be further modified and revised. The proposed framework incorporates a reward model designed to efficiently evaluate the solvability of sub-tasks without requiring actual agent calls. According to the evaluation results, some sub-tasks may be executed by the assigned agents, others may be recognized as inappropriate and be required to perform a replan, while the remaining sub-tasks would be further assessed through the representative works mechanism, which helps determine whether they should be planned in detail or re-described for better aligning the ability of agents. A detector is utilized to identify the missing key information or redundant content in the decomposed sub-tasks, and to provide suggestions to the meta-agent for enhancing the completeness and non-redundancy. Besides, a feedback loop is integrated into the proposed framework to promote ongoing enhancements for the meta-agent.
@@
Extensive experiments are conducted based on a reasoning dataset that needs the collaboration of multiple agents. The comparisons between the proposed agent-oriented planning framework and baseline methods demonstrate the remarkable advancements achieved by the proposed framework. Besides, we conduct an ablation study to confirm the contributions of the different components in the proposed framework, and provide discussions on the potential for further improvements in the performance of agent-oriented planning within multi-agent systems.
@@ 2 Preliminaries
In this study, a multi-agent system involves a series of LLM-empowered agents that possess diverse expertise and tools. These agents are designed to leverage their expertise and tools to resolve specific real-world tasks. For example, a search agent can utilize search engines to retrieve up-to-date information relevant to a given query, while a code agent is capable of generating code and executing the code via a code interpreter. The collaboration among these agents is facilitated by a meta-agent, which serves as the brain of the multi-agent systems. When a user query is submitted to the multi-agent system, the meta-agent sends the relevant requests to the appropriate agents, optimizing for both effectiveness and efficiency. The responses produced by these agents are finally aggregated and synthesized to generate a comprehensive answer to the original user query.
Formally, given a user query $Q$, a meta-agent $\mathcal{P}$ needs to select some appropriate agents from a total of $n$ agents, denoted as $\mathcal{A}=\{\mathcal{A}_{1},...,\mathcal{A}_{n}\}$. Each agent is associated with a description $d$ that outlines its capability. We denote the collection of all agent descriptions as $\mathcal{D}=\{d_{1},...,d_{n}\}$. The user query $Q$ is decomposed into $m$ sub-tasks, with each sub-task assigned to a specific agent based on their descriptions:
@@
After that, each selected agent $i\in[m]$ produces a response to its assigned sub-task, denoted as $r_{i}=\mathcal{A}^{\prime}_{i}(q_{i})$. These responses are then utilized to generate the final answer based on $R=\{r_{1},...,r_{m}\}$ to resolve the original query $Q$. Such a process is called agent-oriented planning.
@@
Secondly, assigning sub-tasks to appropriate agents is non-trivial, as the meta-agent can only rely on the agents’ descriptions to determine task allocation in most cases (Shen et al., 2024). However, a concise and highly generalized description that adequately illustrates an agent’s capabilities may not always be available, leading to suboptimal allocation. For example, if the description of a commonsense agent does not specify the extent of its knowledge base, the meta-agent may struggle to ascertain whether it is suitable to assign the task of querying the melting points of tin and copper to that agent, as shown at the lower left of Figure 1.
@@
Design Principles
We further identify three critical principles that guide the design of our framework for effective and efficient agent-oriented planning in multi-agent systems:
  • Solvability. Each sub-task $q_{i}\ \forall i\in[m]$ should be independently and completely resolvable by at least one single agent within the multi-agent system, ensuring that the response for each sub-task can be reliable. If a sub-task does not satisfy solvability, the meta-agent is expected to take some modifications or further decomposition.
  • Completeness. The array of sub-tasks $\{q_{1},...,q_{m}\}$ should include all necessary information from the original user query $Q$, which ensures that the aggregation of responses of these sub-tasks can effectively yield a comprehensive answer to the user query. If the decomposed sub-tasks fail to satisfy completeness, the decomposition process should be revisited by the meta-agent.
  • Non-Redundancy. The array of sub-tasks $\{q_{1},...,q_{m}\}$ should not include redundant elements, avoiding those task executions that are either irrelevant to resolving $Q$, or duplicated. The principle of non-redundancy promotes that the sub-tasks form a minimal effective set necessary to address the user query, enhancing overall efficiency.
Several examples are provided on the right side in Figure 1 for a better understanding of these design principles. In Section 3, we provide more details of the mechanisms in the proposed framework that are designed to fulfill the aforementioned principles.
@@ 3 Agent-Oriented Planning in Multi-Agent Systems
In this section, we introduce the details of the proposed agent-oriented planning framework. The overall architecture is illustrated in Figure 2.
Figure 2: Overall architecture of the proposed agent-oriented planning framework.
@@ 3.1 Fast Decomposition and Allocation
First of all, following existing studies (Shen et al., 2024; Chen et al., 2023), we provide detailed instructions to the meta-agent to guide it in performing agent-oriented planning. Different from previous studies, the provided instructions incorporate the following requirements: (i) Integration of user query $Q$ and all the agent descriptions $A$: We include both the user query and the descriptions of agents in the instructions, promoting the meta-agent to fully consider the capabilities of each agent and tailor the sub-tasks to align with these capabilities. (ii) Suggestions for assigned agents: We require the meta-agent to provide suggestions for the agents assigned to each decomposed sub-task. Although we tend to separate task decomposition and assignment into two independent tasks performed sequentially, the experimental observations indicate that combining these two tasks enhances the effectiveness of agent-oriented planning. (iii) Structured decomposition of tasks: The decomposed tasks are required to be structured in a sequential manner, specifying any dependencies that may exist between sub-tasks, which ensures that the execution of tasks follows a correct logical order. The adopted prompt can be found in Appendix A.1.
@@
Inspired by recent studies on scaling laws during inference (OpenAI, 2024b), we propose to design mechanisms to guide the meta-agent toward more comprehensive reasoning processes. The results of fast decomposition and allocation should be viewed as intermediate outputs rather than final results, which need to be further evaluated to offer detailed revision suggestions for the meta-agent. More details on this proposed approach are provided in the next section.
@@ 3.2 Is A Sub-Task Completely Resolved?
To tackle this, we propose a reward model designed to provide efficient evaluations of the solvability of sub-tasks, which aims to predict the quality of the agents’ responses to sub-tasks without necessitating actual agent calls.
@@
We utilize a scorer $\mathcal{S}$ that evaluates agents’ responses to sub-tasks, i.e., $\mathcal{S}(q_{i},r_{i,j})=s_{i,j}$, which serves as annotations in the training dataset. The scorer $\mathcal{S}$ provides evaluations from three key aspects: correctness, relevance, and completeness, and can be implemented using powerful LLMs (OpenAI, 2024a) or human annotators. We apply this scoring process across a diverse array of user queries $\mathcal{Q}=\{Q_{k}\}_{k=1}^{K}$ and collect the results to form the training dataset, denoted as $\mathcal{T}=\{(q_{k,i},d_{k,i,j},s_{k,i,j})\ |\ k\in[K],i\in[m_{k}],j\in[l]\}$. More details on the construction of the dataset, such as the adopted prompts, are summarized in Appendix C.
The reward model, parameterized as $\theta$, consists of embedding layers followed by fully connected layers. We obtain embeddings for both the sub-task and the agent’s description separately, then concatenate these embeddings together and feed them into the fully connected layers for further processing. To provide evaluations without making actual agent calls, we design a training objective aimed at minimizing the discrepancies between the model’s predictions and the annotations provided by the scorer, which can be formulated as follows:
@@
With the well-trained reward model, the evaluation of the results produced by fast decomposition and allocation can be summarized as follows. For each suggestion provided by the meta-agent, stating the assignment of sub-task $q_{i}$ to agent $\mathcal{A}_{i}^{\prime}$, the reward model predicts a score denoted as $\mathcal{M}_{\theta}(q_{i},d_{i}^{\prime})$. A sufficiently high predicted score, which implies that $q_{i}$ can be completely resolved by agent $\mathcal{A}_{i}^{\prime}$, leads to the acceptance of this suggestion. Conversely, if the predicted score falls below a predefined threshold, which indicates that the assignment should be revisited by the meta-agent, the reward model provides a set of scores $\hat{s}_{i,j}=\mathcal{M}_{\theta}(q_{i},d_{j})$ for all the agents $\mathcal{A}_{j},j=1,...,n$. From these scores, the meta-agent reports the optimal solution as $j_{max}=\mathrm{argmax}_{j}\,\hat{s}_{i,j}$.
Note that there may be scenarios where the highest score $\hat{s}_{j_{max}}$ is still below the threshold, indicating that the sub-task $q_{i}$ does not satisfy the principle of solvability and that no single agent can resolve it independently. In these cases, the meta-agent $\mathcal{P}$ is required to perform a replan on $q_{i}$. The prompt used for replan is detailed in Appendix A.2.

### v2

@@ Abstract
Through the collaboration of multiple LLM-empowered agents possessing diverse expertise and tools, multi-agent systems achieve impressive progress in solving real-world problems. Given the user queries, the meta-agents, serving as the brain within multi-agent systems, are required to decompose the queries into multiple sub-tasks that can be allocated to suitable agents capable of solving them, so-called agent-oriented planning. In this study, we identify three critical design principles of agent-oriented planning, including solvability, completeness, and non-redundancy, to ensure that each sub-task can be effectively resolved, resulting in satisfactory responses to user queries. These principles further inspire us to propose AOP, a novel framework for agent-oriented planning in multi-agent systems, leveraging a fast task decomposition and allocation process followed by an effective and efficient evaluation via a reward model. According to the evaluation results, the meta-agent is also responsible for promptly making necessary adjustments to sub-tasks and scheduling. Besides, we integrate a feedback loop into AOP to further enhance the effectiveness and robustness of such a problem-solving process. Extensive experiments demonstrate the advancement of AOP in solving real-world problems compared to both single-agent systems and existing planning strategies for multi-agent systems. The source code is available at https://github.com/lalaliat/Agent-Oriented-Planning.
@@ 1 Introduction
In recent years, large language models (LLMs) (Achiam et al., 2023; Chowdhery et al., 2023; Touvron et al., 2023) have achieved impressive breakthroughs in natural language understanding and generation, marking a critical advancement in the exploration of artificial general intelligence (AGI). As the capabilities of LLMs progress, LLM-empowered agents (Qin et al., 2024; Dong et al., 2024; Wang et al., 2024) emerge as key components for integrating expertise and tools to effectively translate these advancements into practical applications. Advancing this paradigm, multi-agent systems (Shen et al., 2023; Gao et al., 2024; Wu et al., 2024), which involve multiple diverse agents, provide great flexibility and adaptability by leveraging and collaborating on the strengths of various agents, promoting comprehensive solutions to complex real-world problems.
@@
Previous studies (Cai et al., 2023; Qian et al., 2024) propose to define suitable standard operating procedures (SOPs) based on the insights and experiences of human professionals for solving specific tasks, such as software development (Hong et al., 2024) and simulating personality traits (Serapio-García et al., 2023). In these scenarios, multiple agents are assigned to execute tasks following the predefined SOPs, resulting in remarkable successes. However, for multi-agent systems designed to tackle a diverse range of complex real-world problems, there is a critical need for a central entity that can automatically generate task-specific operating procedures and coordinate the activities of various agents (Wang et al., 2024).
@@
This central entity, typically referred to as a meta-agent (a.k.a. a controller or planner), carries two primary responsibilities. Firstly, the meta-agent needs to comprehend user queries and decompose them into several sub-tasks, ensuring that each sub-task can be adequately addressed by a single agent. Secondly, the meta-agent is also expected to assign these sub-tasks to the appropriate agents for execution, enabling the solutions of these sub-tasks to collectively provide comprehensive answers to the original user queries.
@@
However, since the meta-agent cannot effectively associate sub-tasks with agents based on the provided agents descriptions (Liu et al., 2024; Hong et al., 2024), the performance of task decomposition and allocations might be sub-optional. To tackle these challenges, we identify three design principles for an agent-oriented planning framework, including solvability, completeness, and non-redundancy. These principles further inspire us to propose AOP, a novel framework for Agent-Oriented Planning in multi-agent systems.
@@
Specifically, to resolve a user query, the meta-agent first performs fast task decomposition and allocation, serving as intermediate results that can be further modified and revised. The proposed AOP incorporates a reward model designed to efficiently evaluate the solvability of sub-tasks without requiring actual agent calls. According to the evaluation results, some sub-tasks may be executed by the assigned agents, others may be deemed inappropriate and require re-planning, while the remaining sub-tasks are further assessed using the representative works mechanism, which helps determine whether they should be planned in detail or re-described for better aligning the ability of agents. Meanwhile, a detector is utilized to identify the missing key information or redundant content in the decomposed sub-tasks, and to provide suggestions to the meta-agent for enhancing the completeness and non-redundancy. Besides, a feedback loop is integrated into AOP to promote ongoing enhancements for the problem-solving process.
@@
Extensive experiments are conducted based on several reasoning datasets that require collaboration among multiple LLM-empowered agents. Comparisons between AOP and baseline methods demonstrate the remarkable advancements achieved by the proposed framework. Furthermore, we conduct an ablation study to show the contributions of different components in AOP, and provide discussions on the potential for further enhancing agent-oriented planning within multi-agent systems.
2 Related Work
Large language models (LLMs) (Brown et al., 2020; Achiam et al., 2023; Chowdhery et al., 2023; Touvron et al., 2023; Le Scao et al., 2023; Zhang et al., 2022) have demonstrated remarkable capabilities in understanding and generating human language (Zhao et al., 2024). To enhance LLMs’ ability to solve complex problems, some research focuses on improving the internal reasoning capabilities of LLMs, typically involving decomposing complex questions into sequential intermediate steps before generating the final responses, as exemplified by Chain-of-Thought (CoT) prompting (Wei et al., 2022) and its variants (Kojima et al., 2022; Wang et al., 2023; Zhou et al., 2023; Li et al., 2023). For further improvements, recent studies propose to adopt multi-agent systems (Wu et al., 2024; Hong et al., 2024; Wu et al., 2025) which leverage the collective intelligence and specialized skills of multiple LLM-empowered agents to solve tasks collaboratively.
Traditional cooperative multi-agent planning methods mainly rely on symbolic methods or reinforcement learning-based methods (He et al., 2016; Yao et al., 2020), with MA-PDDL (Kovacs et al., 2012) being a key example, establishing a standard specification language for multi-agent planning. However, these methods can be limited by their complexity and heavy reliance on human experts (Huang et al., 2024). Recently, the development of LLMs helps to simplify and optimize the planning process through the creation of meta-agents, which serve as controllers that manage the flow of operations, to perform task decomposition (Huang et al., 2024). There are mainly two technical approaches: interleaved decomposition and decomposition-first. Regarding the first kind of approach, React (Yao et al., 2023) generates reasoning and action in an interleaved manner, where reasoning helps the meta-agent update action plans and action enables interaction with agents. HUSKY (Kim et al., 2024) trains an action model to iterate between two stages: generating the next action to take and executing the action using expert models and updating the current solution state. Given that the excessively long trajectories might bring hallucinations in interleaved decomposition, HuggingGPT (Shen et al., 2023) addresses this issue by carefully designing prompts to instruct ChatGPT to decompose tasks, selecting models in HuggingFace and summarizing the generated responses. Chameleon (Lu et al., 2023) prompts the meta-agent with tool descriptions and usage examples to infer a program composed of a sequence of tools to execute for generating the final response.
Although remarkable progress has been made, plans generated solely through prompt-based methods often fall short of meeting the principles of solvability, completeness, and non-redundancy issues that have largely been overlooked in previous studies, which motivates us to propose a novel agent-oriented planning framework for multi-agent systems.
@@ 3 Preliminaries
In this study, a multi-agent system involves a series of LLM-empowered agents for solving real-world problems. These agents are constructed by providing specialized tools and unique system prompts that set their identities and instructions, and are powered by LLMs for query understanding, tool use, and response generation. For example, a search agent can utilize search engines to retrieve up-to-date information relevant to a given query, while a code agent is capable of generating code and executing via a code interpreter. The collaboration among these agents is facilitated by a meta-agent, which serves as the brain of the multi-agent system. When a user query is submitted, the meta-agent sends the relevant requests to the appropriate agents, optimizing for both effectiveness and efficiency. The responses produced by these agents are finally aggregated and synthesized to generate a comprehensive answer to the original user query.
Formally, given a user query $Q$, a meta-agent $\mathcal{P}$ needs to select a set of appropriate agents from a total of $n$ agents, denoted as $\mathcal{A}=\{\mathcal{A}_{1},...,\mathcal{A}_{n}\}$. Each agent is associated with a description $d$ that outlines its capability. We denote the collection of all agent descriptions as $\mathcal{D}=\{d_{1},...,d_{n}\}$. The user query $Q$ can be decomposed by $\mathcal{P}$ into $m$ sub-tasks, with each sub-task assigned to a specific agent based on their descriptions:
@@
After that, each selected agent $i\in[m]$ produces a response to its assigned sub-task, denoted as $r_{i}=\mathcal{A}^{\prime}_{i}(q_{i})$. These responses $\{r_{1},...,r_{m}\}$ are then utilized to generate the final answer to resolve the original query $Q$. Such a process is called agent-oriented planning.
@@
Secondly, assigning sub-tasks to appropriate agents is non-trivial, as the meta-agent can only rely on the agents’ descriptions to determine task allocation in most cases (Shen et al., 2023). However, a concise and highly generalized description that adequately illustrates an agent’s capabilities may not always be available, leading to suboptimal allocation. For example, if the description of a commonsense agent does not specify the extent of its knowledge base, the meta-agent may struggle to ascertain whether it is suitable to assign the sub-task of querying the melting points of tin and copper to that agent, as shown at the lower left of Figure 1.
Design Principles
We further identify three critical principles that guide the design of AOP for effective and efficient agent-oriented planning in multi-agent systems:
  • Solvability. Each sub-task $q_{i},\ \forall i\in[m]$ should be independently and completely resolvable by at least one single agent within the multi-agent system, ensuring that the response for each sub-task can be reliable. If a sub-task does not satisfy solvability, the meta-agent is expected to take some modifications or further decomposition.
  • Completeness. The array of sub-tasks $\{q_{1},...,q_{m}\}$ should include all necessary information from the original user query $Q$, which ensures that the aggregation of responses of these sub-tasks can effectively yield a comprehensive answer to the user query. While a sub-task might include only part of the necessary information, it is not allowable for any particular piece of critical information to be omitted from all sub-tasks.
  • Non-Redundancy. The array of sub-tasks $\{q_{1},...,q_{m}\}$ should not include redundant elements, avoiding those task executions that are either irrelevant to resolving $Q$ or duplicated. The principle of non-redundancy promotes that the sub-tasks form a minimal effective set necessary to address the user query, enhancing overall efficiency.
Several examples are provided on the right side in Figure 1 for a better understanding of these design principles. Note that here we focus on the redundancy between sub-tasks that should be simplified, and believe that adding redundancy among agents is necessary for fault tolerance as discussed in Appendix B, especially considering unexpected changes in agent availability.
While these principles may appear general in nature, we propose an explicit specification and systematic design that follows these principles in the domain of LLM-based multi-agent systems. We hope that these principles, along with the framework we have outlined in the next section, can serve as a good starting point and inspire further research in this field.
@@ 4 Agent-Oriented Planning in Multi-Agent Systems
In this section, we introduce the details of the proposed AOP framework, with the overall architecture illustrated in Figure 2.
Figure 2: Overall architecture of AOP. The framework begins with the meta-agent performing a fast decomposition and allocation of the received user query, resulting in an initial plan. A detector is employed to improve the completeness and eliminate redundancy of the plan, while a reward model provides scores to guide the meta-agent in refining the plan further, which involves operations such as replan, plan-in-detail, etc. The refined plan is sent back to the multi-agent system for generating responses to the user query.
@@ 4.1 Fast Decomposition and Allocation
First of all, following existing studies (Shen et al., 2023; Chen et al., 2024), we provide detailed instructions to the meta-agent to guide it in performing agent-oriented planning. What sets our approach apart is that the provided instructions incorporate the following requirements: (i) Integration of the user query $Q$ and all the agent descriptions $\mathcal{D}$: We include both the user query and the agents’ descriptions in the instructions, promoting the meta-agent to fully consider the capabilities of each agent and tailor the sub-tasks to align with these capabilities. (ii) Suggestions for assigned agents: We require the meta-agent to provide suggestions for the agents assigned to each decomposed sub-task. Although we tend to separate task decomposition and assignment into two independent tasks performed sequentially, the experimental observations indicate that combining these two tasks enhances the effectiveness of agent-oriented planning. (iii) Structured decomposition of tasks: The decomposed tasks are required to be structured in a sequential manner, specifying any dependencies that may exist between sub-tasks, which ensures that the execution of tasks follows a correct logical order. The adopted prompt can be found in Appendix A.1.
@@
Inspired by recent studies on inference time computing (OpenAI, 2024), we propose to design mechanisms to guide the meta-agent toward more comprehensive reasoning processes. The results of fast decomposition and allocation should be viewed as intermediate outputs rather than final results, which need to be further evaluated to offer detailed revision suggestions for the meta-agent. More details on this proposed approach are provided in the next subsection.
@@ 4.2 Is A Sub-Task Completely Resolved?
To tackle this, we introduce a reward model that provides efficient evaluations of the solvability of sub-tasks, which aims to predict the quality of the agents’ responses to sub-tasks without necessitating actual agent calls.
@@
We utilize a scorer $\mathcal{S}$ that evaluates agents’ responses to sub-tasks, i.e., $\mathcal{S}(q_{i},r_{i,j})=s_{i,j}$, which serves as annotations in the training dataset. The scorer $\mathcal{S}$ provides evaluations from three key aspects, including correctness, relevance, and completeness, and can be implemented by using LLMs or employing human annotators. We apply this scoring process across a diverse array of user queries $\mathcal{Q}=\{Q_{k}\}_{k=1}^{K}$ and collect the results to form the training dataset, denoted as $\mathcal{T}=\{(q_{k,i},d_{k,i,j},s_{k,i,j})\ |\ k\in[K],i\in[m_{k}],j\in[l]\}$. More details on the construction of the dataset, such as the adopted prompts, are summarized in Appendix C.
The reward model $\mathcal{M}$, parameterized as $\theta$, consists of embedding layers followed by fully connected layers. We obtain embeddings for both the sub-task and the agent’s description separately, then concatenate these embeddings together and feed them into the fully connected layers for further processing. To provide evaluations without making actual agent calls, we design the training objective aimed at minimizing the discrepancies between the model’s predictions and the annotations provided by the scorer, which can be formulated as follows:
@@
With the well-trained reward model, the evaluation of the results produced by fast decomposition and allocation can be summarized as follows. For each suggestion provided by the meta-agent, stating the assignment of sub-task $q_{i}$ to agent $\mathcal{A}_{i}^{\prime}$, the reward model predicts a score denoted as $\mathcal{M}_{\theta}(q_{i},d_{i}^{\prime})$. A sufficiently high predicted score, which implies that $q_{i}$ can be completely resolved by agent $\mathcal{A}_{i}^{\prime}$, leads to the acceptance of this suggestion. Conversely, if the predicted score falls below a predefined threshold, which indicates that the assignment should be revisited by the meta-agent, the reward model provides a set of scores $\hat{s}_{i,j}=\mathcal{M}_{\theta}(q_{i},d_{j})$ for all the agents $\mathcal{A}_{j},j=1,...,n$. From these scores, the meta-agent reports the optimal selection as $j_{max}=\mathrm{argmax}_{j}\,\hat{s}_{i,j}$.
Note that there may be scenarios where the highest score $\hat{s}_{i,j_{max}}$ is still extremely low, indicating that the sub-task $q_{i}$ does not satisfy the principle of solvability and that no single agent can resolve it independently. In these cases, the meta-agent $\mathcal{P}$ is required to perform a replan on $q_{i}$. The prompt used for replan is detailed in Appendix A.2.